OpenStudy (rock_mit182):

Three consecutive odd integers are such that the square of the third integer is 33 less than the sum of the squares of the first two. One solution is −5​, −3​, and −1. Find three other consecutive odd integers that also satisfy the given conditions. What are the​ integers?

OpenStudy (rock_mit182):

\[X_1 = (2n+1) ; X_2 = (2n+1) +2; X_3 = (2n+1) +4\] \[(X_3)^2 = (X_2 +X_1)^2 -33\]

OpenStudy (rock_mit182):

anyone could check my work ?

OpenStudy (phi):

I think sum of the squares of the first two means \(x_1^2 + x_2^2 \)

OpenStudy (rock_mit182):

oh i see could you help with a hard problem ?

OpenStudy (rock_mit182):

cause i dont know hot translate it into a math equation ..

OpenStudy (rock_mit182):

how*

OpenStudy (rock_mit182):

A 28​-ft by 42​-ft rectangular swimming pool is surrounded by a walkway of uniform width. If the total area of the walkway is 456 ft2​, how wide is the​ walkway? The width of the walkway is _________ feet.

OpenStudy (rock_mit182):

i dont have a clue please help me out

OpenStudy (phi):

can you make that a new post for that question ?

OpenStudy (phi):

for the 3 odd integers, and using a,b, and c a^2 + b^2 = c^2+33 let a= b-2, and c= b+2 (I'm ignoring if b is even or odd... I expect to find an odd answer!) (b-2)^2 + b^2 = (b+2)^2 + 33 expand, simplify, and get \[ b^2 -8b -33=0\] solve for b

OpenStudy (rock_mit182):

ok thank you so much